# Optimizing GPT for Video Understanding: Zero-Shot Performance and Prompt Engineering

### v1

@@ 2 Related Work
In this section, we review topics related to our study, including general and generative multi-modal models.
@@
3 Problem Setup
This section details our methodology for testing the capability of GPT-4o at identifying various Feed Quality (FQ) categories across TikTok. We first provide some relevant background to the problem at hand in Section 3.1, following which we formulate how we test GPT-4o as a vision-language classifier in Section 3.2. We proceed by detailing our results and analysis in Section 4, and conclude by discussing some limitations and future directions in Section 5
Lorem ipsum dolor sit amet, consectetuer adipiscing elit. Ut purus elit, vestibulum ut, placerat ac, adipiscing vitae, felis. Curabitur dictum gravida mauris. Nam arcu libero, nonummy eget, consectetuer id, vulputate a, magna. Donec vehicula augue eu neque. Pellentesque habitant morbi tristique senectus et netus et malesuada fames ac turpis egestas. Mauris ut leo. Cras viverra metus rhoncus sem. Nulla et lectus vestibulum urna fringilla ultrices. Phasellus eu tellus sit amet tortor gravida placerat. Integer sapien est, iaculis in, pretium quis, viverra ac, nunc. Praesent eget sem vel leo ultrices bibendum. Aenean faucibus. Morbi dolor nulla, malesuada eu, pulvinar at, mollis ac, nulla. Curabitur auctor semper nulla. Donec varius orci eget risus. Duis nibh mi, congue eu, accumsan eleifend, sagittis quis, diam. Duis eget orci sit amet orci dignissim rutrum.
Nam dui ligula, fringilla a, euismod sodales, sollicitudin vel, wisi. Morbi auctor lorem non justo. Nam lacus libero, pretium at, lobortis vitae, ultricies et, tellus. Donec aliquet, tortor sed accumsan bibendum, erat ligula aliquet magna, vitae ornare odio metus a mi. Morbi ac orci et nisl hendrerit mollis. Suspendisse ut massa. Cras nec ante. Pellentesque a nulla. Cum sociis natoque penatibus et magnis dis parturient montes, nascetur ridiculus mus. Aliquam tincidunt urna. Nulla ullamcorper vestibulum turpis. Pellentesque cursus luctus mauris.
3.1 Background
@@ 4 Results
Before providing a detailed analysis of the results, we give additional details on the setup of our experiments.

### v2

@@ 2 Related Work
3 Problem and Method
This section details our methodology for testing the capability of GPT-4o at identifying various Feed Quality (FQ) categories across TikTok. We first provide some relevant background to the problem at hand in Section 3.1, following which we formulate how we test GPT-4o as a vision-language classifier in Section 3.2. We proceed by detailing our results and analysis in Section 4, and conclude by discussing some limitations and future directions in Section 5.
3.1 Problem Setup
@@ 5 Conclusion
Limitations: We do not consider finetuning, as our focus is one shot performance. Additionally, we do not list results of other VLM models (e.g. Gemini), but found that initial experiments showed very similar performance between them.